OpenStudy (anonymous):

An auditorium earned $25,000 in sold-out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section

OpenStudy (anonymous):

let F be # of front section tickets ($75 each) and B be # of back section tickets ($50 each). F = 2B (twice as many front section seats as back section) 75F + 50B = 25000 Easy route here would be substitution, as we've already got F in terms of B: 75(2B) + 50B = 25000 150B + 50B = 25000 200B = 25000 B = 125 F = 2B = 2(125) = 250

OpenStudy (anonymous):

Do you get it?